OpenStudy (waheguru):

Students are selling t-shirts to raise money for the recycling club. The profit that they make is given by the expression P = - (21 – 2x)(2 – x).

OpenStudy (waheguru):

In this equation a question is asking what the zeros repersent

OpenStudy (anonymous):

You mean if P=0 or if x=0? If P=0 then it would mean they broke even from selling X amount of shirts, if X is 0 then they did not sell any shirts and P would be negative.

OpenStudy (waheguru):

Can you please explain that more, thats what my teacher said but I did not understand

OpenStudy (waheguru):

Broke even, what does that mean

OpenStudy (anonymous):

So basically, if you want to buy shirts, let us say each shirt cost you $2 to buy, and you buy 10 of them, so you just paid $20 for the shirts, if you sell 0 shirts you lost $20, now lets say you start selling the shirts at $5 each, if you sell 4 shirts at $5 each, you got $20 back, so you spent $20 and made $20, so you just broke even, then any shirt you sell past those 4 you will make a profit on, does that make sense? :)

OpenStudy (waheguru):

Ooh, Thanks soo much

OpenStudy (waheguru):

Also the h = for the vertex is 6.25 but i think its wring and should be 11.5

OpenStudy (waheguru):

Because we add s + r divided by 2

OpenStudy (waheguru):

21 + 2 /2

OpenStudy (anonymous):

Hmm, i'm not sure, if your teacher made up that profit model then it could vary on what they are trying to show.

OpenStudy (waheguru):

How would I find the vertex of this

OpenStudy (anonymous):

What do you mean by vertex? The most amount of money they could make?

OpenStudy (waheguru):

The vertex of the parabola formed by this equation

OpenStudy (waheguru):

@edlevin

OpenStudy (anonymous):

One second let me plot it

OpenStudy (anonymous):

Here is the graph I got: With a max at x=25/4 so P=289/8